OpenStudy (anonymous):

Can someone show me to to find the Domain of Rational expressions on this problem Please? 5m^2+m/7m

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge \frac{5m^2+m}{7m}$$

OpenStudy (anonymous):

5m^2+m/7m m(5m+1)/7m The first thing we have to do is factor the numerator and the denominator. cancel off any common factors. 5m+1/7

OpenStudy (anonymous):

Is that all there is to this problem or am I missing the final step?

OpenStudy (skullpatrol):

You are missing the whole point of the question.

OpenStudy (skullpatrol):

to find the Domain of Rational expressions

OpenStudy (anonymous):

Can you explain this to me?

OpenStudy (skullpatrol):

Sure, the first and most important thing to understand is "division by 0 is undefined."

OpenStudy (skullpatrol):

Do you know why you can never divide by 0?

OpenStudy (anonymous):

because the answer would be zero

OpenStudy (skullpatrol):

If the answer was 0 then it would be OK since 0 is a perfectly valid answer.

OpenStudy (skullpatrol):

-2 + 2 = 0 0*4 = 0

OpenStudy (skullpatrol):

$$ 0\div 6=0$$

OpenStudy (anonymous):

Okay this makes since because the zero is the answer to that problem

OpenStudy (skullpatrol):

Similarly we can have negative numbers as the answers to questions like -6 + 2 = -4 -5*3 = -10 $$-6\div2=-3$$

OpenStudy (anonymous):

Okay if the answer refers to the problem and it ends up being negative.

OpenStudy (skullpatrol):

So with division by 0 the answer is undefined. Because division by 0 is undefined.

OpenStudy (anonymous):

okay

OpenStudy (skullpatrol):

Now look at your question and find out which value of the variable will make it undefined.

OpenStudy (anonymous):

5m^2+m/7m m

OpenStudy (skullpatrol):

which value of the variable will make it undefined.

OpenStudy (anonymous):

I think the "m" that is being added the the 5m^2. I also think the "m" of the 7

OpenStudy (skullpatrol):